MAT 125, CALCULUS A, SUNY AT STONY BROOK, Spring 2004

Text: Single variable calculus: concepts and contexts, James Steward. Syllabus: We will cover approximately Chapters 1,2,3,4. A more detailed syllabus will

be provided. This includes some preliminary material, limits, differentiation and applications.

Classdays. Classes begin on Monday, January 26 and end on Friday, May 7. Holidays: There are no classes during spring break, April 5-11. Lecture times and locations:

day  time  room  instructor
LEC 1 MWF 10:40pm- 11:35am ENGINEERING 143 Kumpel, Paul
R01  M  8:30am-9:25am  OLD CHEM 138  Teh, Jyh-Haur
R02  Th  3:50pm-4:45pm  CHEMISTRY 124  Kim, Young-Deuk
R03  Th  8:20am- 9:15am  PHYSICS P117  An, Daniel
R04  Tu  2:20pm-3:15pm  S B UNION 226  Moraru, Dan
R05  Th  11:20am-12:15am S B UNION 126  Hu, Wenchuan
LEC 2 TuTh 12:50pm-2:10pm ENGINEERING 145 Bishop, Christopher
R06  w  11:45am-12:40pm CHEMISTRY 128  Shu, Yu-Jen
R07  F  11:45am-12:40pm PHYSICS P127  Javaheri, Mohammed
R08  W  9:35am-10:30pm OLD CHEM 138  Shi, Guan-Yu
R09  W  8:30am-9:25am  OLD CHEM 138  Shi, Guan-Yu
R10  F  8:30am-9:25am  OLD CHEM 138  An, Daniel
LEC 3 MW 2:20pm- 3:40pm ENGINEERING 145 Kumpel, Paul
R11  M  9:35am-10:30am  CHEMISTRY 126  Xu, Dezhen
R12  W  10:40am-11:35am CHEMISTRY 124  Javaheri, Mohammed
R13  Tu  3:50pm-4:45pm  CHEMISTRY 123  Thind, Jaimal
R14  Tu  11:20am-12:15pm S B UNION 226  Moraru, Dan
R15  Th  3:50pm-4:45pm  CHEMISTRY 123  Waddington, Travis
ELC 4 TuTh 6:50pm- 8:10pm MATH P-131  Rogers, Luke

There is no recitation for ELC 4. The grader for ELC 4 is Liu, Yuan.

Instructor Office Office hours  email
Bishop  4-112 TuTh 11:30-12:30am  bishop@math.sunysb.edu

 Kumpel  3-120  pkumpel@math.sunysb.edu

Rogers  4-109 W10-12, Th5-6q  lrogers@math.sunysb.edu
An  W10:30-12:30, Th4-5 in MLC  danan@math.sunysb.edu
Hu  Tu5-6 in MLC, 11:20-12:20 in 4-122 wenchuan@math.sunysb.edu

 Javaheri  3-118 W 12:15-2:15  javaheri@math.sunysb.edu
 Kim  Th11-12 in MLC  ydkim@math.sunysb.edu
 Liu  yliu@math.sunysb.edu
 Moraru  moraru@math.sunysb.edu

Shi  2-106 W7-9pm in MLC, Th2-3 in 2-106  guanshi@math.sunysb.edu
Shu  2-114 TuTH 2-3 in MLC, M11-12 in 2-114 yjshu@math.sunysb.edu
Teh  2-120 M2-3 in 2-120, Tu2-4 in MLC  jyhhaur@math.sunysb.edu

 Thind  jthind@math.sunysb.edu
 Waddington  ratatosk@math.sunysb.edu
 Xu  W6-7 in MLC  dezhen@math.sunysb.edu

Homework. A list of problems from the text will be handed out. The recitation instructor will assign, collect and return homework. In general, problems are due the week after the section has been covered in lecture.

Examinations. There will be two midterm tests: Monday, Febuary 23 and Tuesday, March 30. These will be given at night (8:30pm-10:00pm) in a location which will be annouced in class. The final exam is on Monday, May 17 at 11:00-1:30. The location will be announced in class during the last week of the semester. No calculators, notes or books will be allowed during exams. Bring your student ID to all exams.

Early Midterm: There will be an evening exam at 8:30 pm on Wednesday, Febuary 11 in a room to be announced. This is part of the placement process and is given to all students in calculus classes. More information and a practice exam can be found at

http://www.math.sunysb.edu/earlyexam

Grades. Grades are computed using: Recitation 10%, Early exam 10%, Midterms 20% each, Final 40%.

Webpage: Course information is posted at:

http://www.math.sunysb.edu/~bishop/classes/math125.S04/math125.html

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: This is in a new location: S-240A in the basement of the math building. MLC provides help with pre-calculus and calculus material. It is staffed most days and some evenings; an exact schedule will be posted outside the room.

Drop dates: Last day for Add/Drop is January 28. Last day to switch to lower numbered math course is March 19. The last day to withdraw from a course is March 26.